OpenStudy (anonymous):

FAN AND MEDAL! Combine like terms. Like Terms : 1/2 x -1/8x A. 1/8x B. 3/8x C. -1/8x D.-1/16x

OpenStudy (4everaddicted2anime):

First you need to make both of the denominators the same

OpenStudy (4everaddicted2anime):

What did you get when you do that?

OpenStudy (anonymous):

okay so 8/16 and 2/16?

OpenStudy (4everaddicted2anime):

Now you need to subtract the two terms. \[\frac{ 8 }{ 16 }-\frac{ 2 }{ 16 }=?\]

OpenStudy (anonymous):

6/16

OpenStudy (4everaddicted2anime):

Good job. Since that answer isn't one of the choices, you need to simpilfy the term.

OpenStudy (rocketisawesome):

All Wrong

OpenStudy (anonymous):

1/3?

OpenStudy (rocketisawesome):

Not those answers it is B. \[\frac{ 3 }{ 8 }x

OpenStudy (anonymous):

3/8 nvm

OpenStudy (rocketisawesome):

told you

OpenStudy (4everaddicted2anime):

yes. that is your answer

OpenStudy (anonymous):

THANK YOU SO MUCH, I WILL GIVE YOU A MEDAL, AND GIVE You @RocketIsAwesome a fan

OpenStudy (anonymous):

Could you guys help with one more?

OpenStudy (rocketisawesome):

So what are you working on?

OpenStudy (rocketisawesome):

yeah sure

OpenStudy (anonymous):

−22fg − 4f – 9fg − 5f 

OpenStudy (rocketisawesome):

What is a medal for?

OpenStudy (rocketisawesome):

combine like terms?

OpenStudy (rocketisawesome):

-31fg and -9f

OpenStudy (rocketisawesome):

wait no messed up on fg thought it was a 9 it is -26fg and -9f

OpenStudy (rocketisawesome):

nevermind i am sorry

OpenStudy (rocketisawesome):

the first time was correcct

OpenStudy (anonymous):

okay. THANK YOU SO SO MUCH :)

OpenStudy (rocketisawesome):

you gave someone else the medal I think

OpenStudy (anonymous):

Yeah but I gave you a fan :))

OpenStudy (rocketisawesome):

ok

OpenStudy (rocketisawesome):

ill help you with more f you need it